# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v1

@@ 4 Theory
where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the respective orbital degrees of atoms $i$ and $j$, corresponding to the rotation $\mathbf{R}$.
@@
Theorem 2.
The non-linear neural mapping defined as the following is SO(3)-equivariant:
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial(s_{nonlin}(\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)))}{\partial\mathbf{f}}$ (5)
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct product operation, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, and $s_{nonlin}(\cdot)$ represents non-linear neural mappings, $\mathbf{v}$ is the features encoded by $g_{nonlin}(\cdot)$
@@
First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar because the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations. Since applying a non-linear operation to a SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$. It formally holds that:
@@
Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$. Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary.
@@
This proves that $g_{\text{nonlin}}(\cdot)$ is a SO(3)-equivariant non-linear operator: when applying its non-linearity to a SO(3)-equivariant feature $\mathbf{f}$, the output feature $\mathbf{v}$ remains SO(3)-equivariant. ∎
Remark 1.
In Theorem 1, the term $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is equal to the square of the Frobenius norm of the matrix $\mathbf{Q}$, i.e., $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\|\mathbf{Q}\|_{F}^{2}=\sum_{m,n}|\mathbf{Q}_{mn}|^{2}$ The invariance of Frobenius norm under unitary transformations reflects the profound connections between modulus and invariance in physics.
Remark 2.
In Theorem 2, $g_{\text{nonlin}}(\cdot)$ can handle equivariant features of any degree and the returned feature $\mathbf{v}$ has the same degree $l$ as the input feature $\mathbf{f}$. $\mathbf{v}$ is endowed with non-linear expressiveness with the effect of $s_{nonlin}(\cdot)$, which can be any differentiable non-linear network module, including non-linear activation functions and normalization operations. The gradient operation $\frac{\partial z}{\partial\mathbf{f}}$ in $g_{\text{nonlin}}(\cdot)$ is used during the forward inference process to promote feature encoding, expanding the usage and significance of gradients within the neural network paradigm. It is worth noting that, although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance. Literature [7] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts.
Remark 3.
$z$ in Equation (6) is the bridge connecting Theorem 1 and Theorem 2 in implementing a SO(3)-equivariant non-linear representation learning approach. Since $z$ is an SO(3)-invariant feature, it can be learned by $s_{\text{nonlin}}(\cdot)$, i.e., the neural mechanisms with non-linear expressive capability, through supervised learning incorporated with the SO(3)-invariant label $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$. Following this, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as proved above. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations.
Based on the new theory, we present a general method for learning non-linear representations that are equivariant to SO(3), as shown in Figure 1. The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates an SO(3)-invariant trace quantity introduced in Theorem 1 as a supervisory signal for learning SO(3)-invariant features; at the representation level, it induces non-linear SO(3)-equivariant features from non-linear SO(3)-invariant features through gradient-based operator introduced in Theorem 2. For details of the method, please refer to Section 8.
5 Application to the Hamiltonian Prediction Tasks
The tasks of predicting electronic-structure Hamiltonian are chosen as the application study of our theoretical and methodological framework, on two motivations. On one hand, the Hamiltonian occupies a central position in quantum mechanics. Form the Hamiltonian, key physical quantities such as charge density, band structure, and Berry phase can be calculated, bringing deeply understanding of the electrical, optical, magnetic, and transport properties of atomic systems. On the other hand, the intrinsic symmetry with respect to the SO(3) group as well as the high-dimensional complexity (as shown in Figure 2) of the Hamiltonian provide a challenging touchstone of whether a deep learning method possesses both SO(3)-equivariance and strong expressive power.
Figure 2: Illustration of the high-dimensional complexity of the Hamiltonian. In an atomic system, a Hamiltonian matrix is defined between each pair of atoms within each other’s cutoff radius, constituting a large cluster of Hamiltonian matrices for all atom pairs. Besides, each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here. Thus, for a Hamiltonian prediction model, the space of the regression target is typically huge.
Experiments are conducted on six benchmark databases of atomic systems that have Hamiltonian annotations, all of which are established and released by previous literature [6, 7]. These databases include:
• Monolayer Graphene (MG). This structure consists of a single layer of carbon atoms in a hexagonal lattice. It is known for its exceptional thermal conductivity, electrical properties, and mechanical strength. The potential applications of monolayer graphene extend across electronics, energy storage devices, and composite materials due to its unique properties. In the benchmark of MG, the sizes of the training set, validation set, and testing set are 270, 90, and 90 respectively.
• Monolayer MoS2 (MM), which is a single layer of molybdenum disulfide and part of the transition metal dichalcogenides (TMDCs) group. This kind of structures is a semiconductor with a direct bandgap, making it suitable for optoelectronics, photovoltaics, and semiconducting components. It also shows promise in applications requiring high sensitivity to chemicals, making it valuable for sensors. In the benchmark of MM, the sizes of the training set, validation set, and testing set are 300, 100, and 100 respectively.
• Bilayer Graphene (BG), which consists of two layers of graphene stacked together. This kind of structures exhibits unique electronic properties such as tunable bandgap and enhanced conductivity when subjected to an electric field. Bilayer Graphene is being explored for use in advanced electronic devices, such as fast transistors and efficient batteries. In the benchmark database of BG, the sizes of the training set and validation set are 180 and 60 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 60 and 9 respectively.
• Bilayer Bismuthene (BB), which consists of two layers of bismuth atoms arranged in a honeycomb-like structure. This kind of materials shows potential for topological insulating properties and has significant implications for quantum computing and spintronic devices due to its strong spin-orbit coupling. In the benchmark of BB, the sizes of the training set and validation set are 231 and 113 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 113 and 4 respectively.
• Bilayer Bi2Te3 (BT), which is a compound of bismuth and tellurium, known for its thermoelectric properties. This kind of structures is particularly valuable in thermoelectric devices for power generation and refrigeration technologies where waste heat can be converted into electrical energy. In the benchmark of BT, the sizes of the training set and validation set are 204 and 38 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 12 and 2 respectively.
• Bilayer Bi2Se3 (BS), another bismuth-tellurium compound, similar to Bi2Te3 but with selenium replacing tellurium. It also exhibits excellent thermoelectric properties and potential for topological insulator applications. This makes it suitable for energy harvesting applications and advanced electronic devices based on topological insulator technology. In the benchmark of BS, the sizes of the training set and validation set are 231 and 113 respectively, all of them are non-twisted samples; in the testing set, the numbers of non-twisted and twisted samples are 113 and 2 respectively.
These databases are both diverse and representative. They encompass both strong chemical bonds between single-layer atoms and weaker van der Waals interactions between bilayers, and also cover varying degrees of Spin-Orbit Coupling (SOC) effects: structures like BT, BB, and BS exhibit strong SOC effects, while others have weak SOC. These atomic structures are highly valuable in the field of information and energy science and technology. Yet, accurately predicting their Hamiltonians poses significant challenges due to structural deformations from thermal movements and inter-layer twists, as depicted in Fig. 3, where the twisted structures have garnered considerable attention as a research focus because of their potential to exhibit novel electrical properties and quantum topological behaviors [24, 25]. During twisting, the relative rotation between atoms and their coordinate system introduces SO(3)-equivariant effects, while changes in orientation between two layers alter the vdW interactions. These dynamic conditions challenge both the equivariance and expressive capabilities of the regression model. In our experiments, predicting the Hamiltonian of twisted samples are even challenging, as we strictly adhere to the official division of training, validation, and testing sets as defined upon the dataset’s release [^1] https://github.com/Xiaoxun-Gong/DeepH-E3. In this division, twisted samples are only included in the testing set and are absent from the training set. This setup forces the regression model to solely rely on its generalization ability and the inherent SO(3)-equivariance to extrapolate the twisting conditions.
In summary, these benchmarks present scientific challenges for a regression model, pushing the boundaries of modeling capabilities in predicting complex quantum behaviors under dynamic conditions.
Figure 3: Illustration of challenging samples in the experimental databases, showing structural distortions of atomic systems due to thermal movements and twists between the two layers, which require a regression model to possess both strict SO(3)-equivariance and strong expressive capacity.
6 Results and Analysis
To validate the effectiveness of the proposed method, we compare results from the four setups:
• Baseline model (DeepHE3 [7]): We select DeepHE3 as our baseline model since it represents the state-of-the-art (SOTA) among strictly SO(3)-equivariant approaches for predicting electronic-structure Hamiltonians. This choice facilitates further breakthroughs based on the SOTA foundation. DeepHE3 is trained using the mean squared error (MSE) loss of the Hamiltonian. In the subsequent experimental setups, DeepHE3 will serve as a base encoder, providing basic SO(3)-equivariant features that are handled by our method as more expressive representations.
@@ 6 Results and Analysis
• DeepHE3+Trace: This experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepHE3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include mechanisms for encoding non-linear SO(3)-equivariant features via the gradient-induced operator; instead, it directly uses the features outputted by DeepHE3 for Hamiltonian regression.
@@
• DeepHE3+Grad: This setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term. Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the base encoders of DeepHE3 for Hamiltonian prediction. As for ablation study, this setup continues to use the single-task training pipeline of DeepHE3, predicting only the Hamiltonian without joint supervision training through the trace label.
@@
• DeepHE3+TraceGrad. This setup corresponds to a complete implementation of our framework. Extending beyond the baseline architecture and pipeline of DeepHE3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage non-linear SO(3)-invariant features induced by the gradient operator to yield non-linear SO(3)-equivariant features for Hamiltonian prediction.
The performance of interest for the electronic-structure Hamiltonian prediction task are SO(3)-equivariance and numerical accuracy. Our method has already been theoretically proven to strictly obey SO(3)-equivariance in Section 4; therefore, we just focus on evaluating numerical accuracy here. Four metrics designed by Yin et al. [16] are adopted, i.e., $MAE_{all}$, which calculates the Mean Absolute Error (MAE) averaged across all elements of the Hamiltonian matrices from all testing samples in the current database; $MAE^{cha}_{sample}$, which calculates the MAE on the most challenging testing samples where the baseline model (DeepHE3) performs the worst in the current database; $MAE_{block}$, which calculates the MAE for each basic block of the Hamiltonian matrix, averaged across all testing samples in the current database; $MAE^{cha}_{block}$, which corresponds to the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE_{block}$) in the current dataset. These metrics comprehensively reflect the accuracy performance, considering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices.
We collect results from $10$ independently repeated experiments on each database and list them in Tables 1 and 2, including the values of $MAE_{all}$, $MAE^{cha}_{sample}$, and $MAE^{cha}_{block}$, for monolayer and bilayer structures, respectively. Considering the large volume of statistics on $MAE_{block}$ for all blocks, we just display the complete results in Figures 4 and 5 with the form of bar charts for DeepHE3 and DeepHE3+TraceGrad. The results for DeepHE3 listed in these tables and figures are directly taken from the reproduction of Yin et al. [16], a paper which is in fact previous work by our own team. Since the experimental conditions for DeepHE3 are exactly the same between the previous and current work of ours, we directly list the results from the previous paper here. As pointed out in [16], the reproduction was based on DeepHE3’s open-source project11footnotemark: 1, which is newer than its original publication, and the performance was slightly better than the original one. As we use a fixed random seed ($42$) for all of the experiments, we find that the standard deviation of each metric is very small (all less than $0.007$ meV) across the $10$ repeated experiments. Since the fluctuation is negligible, in the following tables and figures, we just present the mean values of $MAE_{all}$, $MAE^{cha}_{sample}$, and $MAE^{cha}_{block}$ across the repeated experiments, without listing the standard deviations of each entry to save space.
Table 1: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics (/meV) for the four experimental setups on the databases of monolayer structures.
@@
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
DeepHE3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
DeepHE3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
DeepHE3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
Table 2: Comparison on the $MAE_{all}$, $MAE^{cha}_{sample}$ and $MAE^{cha}_{block}$ metrics (/meV) for the four experimental setups on the databases of bilayer structures, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
@@
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
DeepHE3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
DeepHE3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
DeepHE3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
@@
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
DeepHE3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
DeepHE3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
DeepHE3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
@@
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
DeepHE3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
DeepHE3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
DeepHE3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
@@
$MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$ $MAE_{all}$ $MAE^{cha}_{sample}$ $MAE^{cha}_{block}$
DeepHE3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
DeepHE3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
DeepHE3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
DeepHE3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674
Figure 4: $MAE_{\text{block}}$ on each basic block of the Hamiltonian matrices of monolayer structures.
Combining the results from Table 1, Table 2, Figure 4 and Figure 5, we find that the proposed TraceGrad method significantly enhances the accuracy performance of the baseline method DeepHE3, both on average and for challenging samples or blocks, both on the non-twisted samples and the twisted samples marking strong difficulty for the expressive power and generalization ability of regression models as analyzed in Section 5. Specifically, TraceGrad lowers down the $MAE_{all}$ of DeepHE3 with a relative ratio of $30.28\%$, $29.80\%$, $25.19\%$, $25.00\%$, $17.52\%$, $17.95\%$, $34.00\%$, $11.55\%$, $24.43\%$, and $21.35\%$, on MG, MM, BGnt, BGt, BBnt, BBt, BTnt, BTt, BSnt, and BSt, respectively; and lowering down $MAE_{cha}$ by $28.01\%$, $28.22\%$, $28.70\%$, $13.29\%$, $15.79\%$, $16.45\%$, $35.00\%$, $11.18\%$, $21.70\%$, and $22.56\%$, on the respective datasets; and achieving better results in the vast majority of basic blocks. Particularly, for the blocks where DeepHE3 performs the worst, TraceGrad reduces the $MAE$ ($MAE^{cha}_{block}$) by $37.01\%$, $26.75\%$, $33.23\%$, $33.33\%$, $28.98\%$, $46.48\%$, $48.23\%$, $3.37\%$, $25.72\%$, and $22.97\%$, on the respective datasets, from which the $MAE^{cha}_{block}$ in MM, BGnt and BTnt are lowered down to within $1meV$; the $MAE^{cha}_{block}$ in BBt is decreased by nearly half to approach $1meV$. This indicates that the proposed representation learning framework, while maintaining strict SO(3)-equivariance, significantly enhances the expressiveness of equivariant representations, thereby showcasing immense advantages in capturing the complexity of physical systems and refreshing the SOTA accuracy on the six benchmark databases.
Figure 5: $MAE_{\text{block}}$ on each basic block of the Hamiltonian matrix in non-twisted (marked by superscripts $nt$) and twisted (marked by superscripts $t$) bilayer structures.

### v3

@@ B.1 Proof of Theorem 1
where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l_{p}$ and $l_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
@@
B.2 Proof of Theorem 2
@@
First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar because the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations. Since applying a non-linear operation to an SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$. It formally holds that:
@@
Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$. Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary  [^3] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, because the output space may be complex while the internal representations of neural networks are typically real. Although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance. Literature [8] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts..
@@
This proves that $g_{\text{nonlin}}(\cdot)$ is an SO(3)-equivariant non-linear operator: when applying its non-linearity to a SO(3)-equivariant feature $\mathbf{f}$, the output feature $\mathbf{v}$ remains SO(3)-equivariant. ∎
Appendix C Information of Experimental Databases
In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH series [6, 8], and the two databases from the QH9 series [19], listed in Table 5 and Table 6, respectively. Additionally, we visualize two types of challenging testing samples: samples with non-rigid deformation from thermal motions, as well as the bilayer samples with interlayer twists, which are shown in Fig. 4 and Fig. 5, respectively.
Table 5: Statistic information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series [6, 8]. SOC: effects of Spin-Orbit Coupling. $m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset. $nt$: non-twisted samples. $t$: twisted samples.
Statistic Types MG MM BG BB BT BS
Elements C Mo, S C Bi Bi, Te Bi, Se
SOC weak weak weak strong strong strong
Training ($nt$) $m$ 270 300 180 231 204 231
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Validation ($nt$) $m$ 90 100 60 113 38 113
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Testing ($nt$) $m$ 90 100 60 113 12 113
$a_{max}$ 72 75 64 36 90 90
$a_{min}$ 72 75 64 36 90 90
Testing ($t$) $m$ 9 4 2 2
$a_{max}$ 1084 244 130 190
$a_{min}$ 28 28 70 70
Table 6: Basic statistics of the two benchmark databases, QH9-Stable ($QS$) and QH9-Dynamic ($QD$), from the QH9 benchmark series [19]. For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and test sets, ensuring that the atom number of samples do not overlap across the three subsets. For the $QD$ database, the ’mol’ strategy from the official settings is employed to split the training, validation, and test sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets. These split strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories. $m$: number of samples in the current dataset. $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset.
Statistic Types QS QD
Elements C, H, O, N, F C, H, O, N, F
Training $m$ 104,001 79,900
$a_{max}$ 20 19
$a_{min}$ 3 10
Validation $m$ 17,495 9,900
$a_{max}$ 22 19
$a_{min}$ 21 10
Testing $m$ 9,335 10,100
$a_{max}$ 29 19
$a_{min}$ 23 10
Figure 4: Visualization of testing samples exhibiting non-rigid deformations due to thermal motions
Figure 5: Visualization of testing samples with interlayer twists.
Appendix D Training Configurations
$\lambda$ in the training loss function is set as $0.3$ according to parameter selection on the validation sets. The initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$ and $BS$ databases are set as $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively, the training batch size for them is set as $1$ and the optimizer is selected as Adam; the initial learning rate for experiments on the $QS$ and $QD$ databases is set as $0.0005$, the training batch size is set as $16$, and the optimizer is selected as AdamW.
Appendix E Visualization of Block-level MAE Statistics
The MAE metrics (denoted as $MAE^{H}_{block}$) for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, across different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig. 6 and 7. Fig. 6 presents the results for monolayer structures, while Fig. 7 focuses on bilayer structures. From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, i.e., DeepH-E3, on most of the blocks of the Hamiltonian matrix, especially on those blocks where DeepH-E3 performs poorly.
Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the $MG$ and $MM$ databases.
Figure 7: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) structures of $BG$, $BB$, $BT$ and $BS$.
Appendix F Ablation Study
We conduct fine-grained ablation study on the six databases from DeepH benchmark series, comparing results from the four setups:
• DeepH-E3 [8]: the baseline model.
@@ Appendix F Ablation Study
• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include the gradient-based mechanism for encoding SO(3)-equivariant non-linear features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression.
@@
• DeepH-E3+Grad: this setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term. Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the baseline encoders of DeepH-E3 for Hamiltonian prediction. As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, predicting only the Hamiltonian without joint supervision training through the trace label.
@@
• DeepH-E3+TraceGrad: this is a complete implementation of our framework extending beyond of the architecture and training pipeline of DeepH-E3, at the label level, we introduce the trace quantity to guide the learning of SO(3)-invariant features; Meanwhile, at the representation level, we leverage the gradient operator to yield SO(3)-equivariant non-linear features for Hamiltonian prediction.
Table 7: Ablation study MAE results (meV) on the $MG$ and $MM$ databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.251 0.357 0.362 0.406 0.574 1.103
DeepH-E3+Trace 0.230 0.344 0.348 0.378 0.537 1.091
DeepH-E3+Grad 0.185 0.269 0.258 0.308 0.453 0.924
DeepH-E3+TraceGrad 0.175 0.257 0.228 0.285 0.412 0.808
Table 8: Ablation study MAE results (meV) on $BG$, $BB$, $BT$and $BS$. The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.389 0.453 0.644 0.264 0.429 0.609
DeepH-E3+Trace 0.362 0.417 0.593 0.251 0.401 0.480
DeepH-E3+Grad 0.320 0.356 0.511 0.222 0.389 0.446
DeepH-E3+TraceGrad 0.291 0.323 0.430 0.198 0.372 0.406
@@
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.274 0.304 1.042 0.468 0.602 2.399
DeepH-E3+Trace 0.259 0.285 0.928 0.429 0.570 1.782
DeepH-E3+Grad 0.243 0.272 0.824 0.406 0.542 1.431
DeepH-E3+TraceGrad 0.226 0.256 0.740 0.384 0.503 1.284
@@
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.447 0.480 1.387 0.831 0.850 4.572
DeepH-E3+Trace 0.406 0.462 1.239 0.784 0.812 4.520
DeepH-E3+Grad 0.342 0.365 0.750 0.742 0.786 4.463
DeepH-E3+TraceGrad 0.295 0.312 0.718 0.735 0.755 4.418
@@
$MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$ $MAE_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-E3 (Baseline) 0.397 0.424 0.867 0.370 0.390 0.875
DeepH-E3+Trace 0.382 0.397 0.843 0.351 0.367 0.838
DeepH-E3+Grad 0.343 0.365 0.696 0.324 0.339 0.746
DeepH-E3+TraceGrad 0.300 0.332 0.644 0.291 0.302 0.674